# Contact non-squeezing in lens spaces

### v1

@@ Abstract.
We describe some contact non-squeezing phenomena in lens spaces by defining and computing a contact capacity. This contact capacity comes from the spectral selectors constructed by Allais, Sandon and the author by the means of generating functions and Givental’s non-linear Maslov index [4]. We also discuss a potential generalization of these non-squeezing phenomena for orderable closed prequantizations.
@@ 1.1. Historical backgroud and main result
For contact geometry, considered sometimes as the odd-dimensional analogue of symplectic geometry, the story is somewhat different. In this geometry, the volume is not an invariant anymore and, depending on the contact manifold considered, global invariants might not exist at all. For example, in the standard contact Euclidean space, which is the local model of contact geometry, any bounded domain can be contactly squeezed inside any arbitrary small neighborhood. Therefore, in a general contact manifold, non-squeezing phenomena cannot be expected at small scale. Nonetheless, 20 years after Gromov’s result Eliashberg, Kim and Polterovich [14] overcame this additional difficulty and proved a non-squeezing theorem at large scale in the prequantization of the standard symplectic Euclidean space. Unlike symplectic geometry, generalization of non-squeezing phenomena to other contact manifolds remain however rare. To the knowledge of the author, such phenomena have been detected only in prequantization of Liouville manifolds ([1, 6, 2, 32, 16, 8, 18]), some exotic contact spheres [35] and the unitary cotangent of the torus endowed with its canonical contact structure (see for example [13, Thm 2.4.1]).
In this paper we show that interesting non-squeezing phenomena happen also in lens spaces endowed with their standard contact structure.
Let us fix $n\in\mathbb{N}_{>0}$ a positive integer. Endow the Euclidean space $\mathbb{R}^{2n+2}$ with the coordinate functions $((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))$ and let $\mathbb{S}^{2n+1}:=\left\{\sum\limits_{j=1}^{n+1}x_{j}^{2}+y_{j}^{2}=1\right\}$ be the standard Euclidean sphere of radius $1$ centered at 00. Identifying $\mathbb{R}^{2n+2}$ with $\mathbb{C}^{n+1}$ via $\Phi:((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))\mapsto(x_{1}+iy_{1},\cdots,x_{n+1}+iy_{n+1})$ we get an action of $\mathbb{S}^{1}:=\{e^{i\theta}\ |\ \theta\in\mathbb{R}\}\subset\mathbb{C}$ on $\mathbb{S}^{2n+1}\subset\mathbb{C}^{n+1}$ given by multiplication. For all integer $k\geq 1$, the lens space $L_{k}^{2n+1}$ is the quotient space $\mathbb{S}^{2n+1}/\mathbb{Z}_{k}$ where $\mathbb{Z}_{k}:=\mathbb{Z}/k\mathbb{Z}\simeq\{e^{\frac{2i\pi m}{k}}\ |\ m\in[0,k-1]\cap\mathbb{N}\}$ is identified with the $k$th roots of unity. We denote by $\pi_{1}^{k}:\mathbb{S}^{2n+1}\to L_{k}^{2n+1}$ the covering map. It carries a natural contact form $\alpha_{k}$ defined as follow: $(\pi^{k}_{1})^{*}\alpha_{k}=\alpha_{1}$ where $\alpha_{1}=i_{\mathbb{S}^{2n+1}}^{*}\left(\sum\limits_{j=1}^{n+1}(x_{j}dy_{j}-y_{j}dx_{j})\right)$ and $i_{\mathbb{S}^{2n+1}}:\mathbb{S}^{2n+1}\hookrightarrow\mathbb{R}^{2n+2}$ is the inclusion.
@@
    In [26, Thm 1 and Prop 2.11] [^1] The conventions used in [26] are different: they consider the symplectic form $\widetilde{\omega}:=\frac{1}{2\pi}{\omega}$ and show that there exists an embedding of $(\widetilde{\mathbb{B}^{2n}}(1),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\widetilde{\omega})$, where $\widetilde{\mathbb{B}^{2n}}(a):=\mathbb{B}^{2n}(\sqrt{\frac{a}{\pi}})$ Karshon and Tolman construct a symplectic embedding $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$, for $R=\sqrt{2}$, and show that it is optimal in the sense that such symplectic embeddings do not exist for $R>\sqrt{2}$ (see also [28, Thm 1]). Considering this optimal embedding, one can see that non-squeezing phenomena appear for all lens spaces, even for the real projective space $(\mathbb{R}\mathrm{P}^{2n+1},\xi):=(L_{2}^{2n+1},\xi_{2})$, since $\pi\left(\sqrt{2}\right)^{2}>\frac{2\pi}{2}$.
@@
We will prove Theorem 1.1 by constructing a contact capacity defined on open subsets of $(L_{k}^{2n+1},\xi_{k})$ and by estimating this contact capacity for the lift of Darboux balls. This contact capacity comes from the spectral selectors defined on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ in [4].
@@ 1.2. Discussion
It is interesting to note that the $\alpha_{k}$-Reeb flow in $L_{k}^{2n+1}$ is periodic with period $\frac{2\pi}{k}$ for all integers $k\geq 1$. Indeed, the periodicity of the Reeb flow often plays a crucial role in the existence of non trivial global invariants and this article is no exception to this fact (see for example Remark 2.4 and Lemma 6.2). Furthermore, the global invariants one can deduce from Theorem 1.1 naturally take value in the discrete set consisting in multiples of this period. Similarly, in the contact manifold $\left(\mathbb{R}^{2n}\times\mathbb{R}/\mathbb{Z},\alpha_{0}:=\mathrm{d}z-\frac{1}{2}\sum\limits_{i=1}^{n}(x_{i}dy_{i}-y_{i}dx_{i})\right)$ where the first contact non-squeezing phenomena was described by Eliashberg, Kim and Polterovich in [14], the $\alpha_{0}$-Reeb flow is $1$-periodic, and they show that it is impossible to contactly squeeze $\mathbb{B}^{2n}(A)\times\mathbb{R}/\mathbb{Z}$ inside $\mathbb{B}^{2n}(a)\times\mathbb{R}/\mathbb{Z}$ if there exists an integer $j\in\mathbb{N}_{>0}$ so that $A>j\geq a$.
@@
It seems that another important property that a contact manifold should have so that one could expect meaningful invariant measurements is to be orderable (see Section 6 for a definition). For $n\in\mathbb{N}_{>0}$ and $k\geq 2$, the spheres $(\mathbb{S}^{2n+1},\ker\alpha_{1})$ and the lens spaces $(L_{k}^{2n+1},\ker\alpha_{k})$ are perfect representative cases since the sphere is not orderable [14] while the lens space is [21], and, as we have just mentionned, interesting contact measurements exist for the latter but not for the former. In the same direction, a more general observation can be made for bi-invariant metrics on contactomorphisms groups. On the one hand, to the knowledge of the author, the only examples in the litterature of closed contact manifolds $(M,\xi)$ admitting unbounded bi-invariant metrics on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ are the contact manifolds $(M,\xi)$ that are orderable and that admit a periodic Reeb flow [17, 10, 21, 4]. On the other hand, for a large family of closed non-orderable contact manifolds $(M,\xi)$ Courte and Massot show that $\mathrm{Cont}_{0}(M,\xi)$ and $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ have to be bounded [11].
That being so, we discuss in Section 6 a potential generalization of Theorem 1.1 to orderable prequantizations of closed symplectic manifolds which are particular cases of orderable closed contact manifold with periodic Reeb flow.
Actually, while convenient but not sufficient, the existence of a periodic Reeb flow is not even a necessary condition to have contact global invariants. For example, the unitary cotangent of the torus $\mathbb{P}^{+}T^{*}\mathbb{T}^{n}$ endowed with its canonical contact structure $\xi_{\mathrm{can}}$ does not admit any periodic Reeb flow when $n>1$ (see for example [7]). But, as briefly mentionned above, one can use an interesting contact global invariant for domains in $(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$, which is called the shape333unlike the other invariant measurements discussed so far, the shape can be used to construct “non-discrete” measurements, to deduce non-squeezing phenomena and other rigidity results [13, 6, 29, 15]. It would therefore be interesting to investigate whether the groups $\widetilde{\mathrm{Cont}_{0}}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ and $\ \mathrm{Cont}_{0}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ are bounded or not (see also [3, Prop. 4.14, Rem. 4.15]).
@@
In Section 2 we recall briefly the construction and main properties of the spectral selectors constructed in [4]. In Section 3 we define the above mentionned contact capacity for open domains of $(L_{k}^{2n+1},\xi_{k})$ coming from the spectral selectors. In this same Section 3 we state - without prooving - some properties of this contact capacity (see Theorem 3.1) which allow us to give the proof of Theorem 1.1. In Section 4 we show a energy-capacity inequality satisfied by the contact capacity. In Section 5 we prove Theorem 3.1 by estimating and computing the contact capacity of domains that are lifts of Darboux balls using the energy-capacity inequality of Section 4 and some explicit constructions. Finally, we discuss in Section 6 on a possible generalization of this result to orderable closed prequantizations.

### v2

# Contact non-squeezing in various closed prequantizations

@@ Abstract.
We describe some contact non-squeezing phenomena, first in lens spaces then in strongly orderable closed prequantizations in the sense of Liu [30]. To detect and quantify these non-squeezing phenomena we define and compute two different contact capacities. In the former case the contact capacity comes from the spectral selectors constructed by Allais, Sandon and the author by the means of generating functions and Givental’s non-linear Maslov index [4]. In the latter case the contact capacity comes from the order spectral selectors constructed by Allais and the author [3].
@@ 1. Introduction
For contact geometry, considered sometimes as the odd-dimensional analogue of symplectic geometry, the story is somewhat different. In this geometry, the volume is not an invariant anymore and, depending on the contact manifold considered, global invariants might not exist at all. For example, in the standard contact Euclidean space, which is the local model of contact geometry, any bounded domain can be contactly squeezed inside any arbitrary small neighborhood. Therefore, in a general contact manifold, non-squeezing phenomena cannot be expected at small scale. Nonetheless, 20 years after Gromov’s result Eliashberg, Kim and Polterovich [15] overcame this additional difficulty and proved a non-squeezing theorem at large scale in the prequantization of the standard symplectic Euclidean space. Unlike symplectic geometry, generalization of non-squeezing phenomena to other contact manifolds remain however rare. To the knowledge of the author, such phenomena have been detected only in prequantization of Liouville manifolds ([1, 5, 2, 36, 17, 9, 19]), some exotic contact spheres [38] and the unitary cotangent of the torus endowed with its canonical contact structure (see for example [14, Thm 2.4.1]).
In this paper we show that interesting non-squeezing phenomena happen also in lens spaces endowed with their standard contact structure and in strongly orderable closed prequantizations.
1.1. Non-squeezing in contact lens spaces
Let us fix $n\in\mathbb{N}_{>0}$ a positive integer. Endow the Euclidean space $\mathbb{R}^{2n+2}$ with the coordinate functions $((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))$ and let $\mathbb{S}^{2n+1}:=\left\{\sum\limits_{j=1}^{n+1}x_{j}^{2}+y_{j}^{2}=1\right\}$ be the standard Euclidean sphere of radius $1$ centered at 00. Identifying $\mathbb{R}^{2n+2}$ with $\mathbb{C}^{n+1}$ via $\Phi:((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))\mapsto(x_{1}+iy_{1},\cdots,x_{n+1}+iy_{n+1})$ we get an action of $\mathbb{S}^{1}:=\{e^{i\theta}\ |\ \theta\in\mathbb{R}\}\subset\mathbb{C}$ on $\mathbb{S}^{2n+1}\subset\mathbb{C}^{n+1}$ given by multiplication. For all integer $k\geq 1$, the lens space $L_{k}^{2n+1}$ is the quotient space $\mathbb{S}^{2n+1}/\mathbb{Z}_{k}$ where $\mathbb{Z}_{k}:=\mathbb{Z}/k\mathbb{Z}\simeq\{e^{\frac{2i\pi m}{k}}\ |\ m\in[0,k-1]\cap\mathbb{N}\}$ is identified with the $k$th roots of unity. It carries a contact distribution $\xi_{k}$ given by the kernel of the contact form $\alpha_{k}$ defined as the pullback, by the covering map $\mathbb{S}^{2n+1}\to L_{k}^{2n+1}$, of $\alpha_{1}:=i_{\mathbb{S}^{2n+1}}^{*}\left(\sum\limits_{j=1}^{n+1}(x_{j}dy_{j}-y_{j}dx_{j})\right)$ where $i_{\mathbb{S}^{2n+1}}:\mathbb{S}^{2n+1}\hookrightarrow\mathbb{R}^{2n+2}$ is the inclusion.
@@ 1.1. Non-squeezing in contact lens spaces
    In [27, Thm 1 and Prop 2.11] [^1] The conventions used in [27] are different: they consider the symplectic form $\widetilde{\omega}:=\frac{1}{2\pi}{\omega}$ and show that there exists an embedding of $(\widetilde{\mathbb{B}^{2n}}(1),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\widetilde{\omega})$, where $\widetilde{\mathbb{B}^{2n}}(a):=\mathbb{B}^{2n}(\sqrt{\frac{a}{\pi}})$ Karshon and Tolman construct a symplectic embedding $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$, for $R=\sqrt{2}$, and show that it is optimal in the sense that such symplectic embeddings do not exist for $R>\sqrt{2}$ (see also [31, Thm 1]). Considering this optimal embedding, one can see that non-squeezing phenomena appear for all lens spaces, even for the real projective space $(\mathbb{R}\mathrm{P}^{2n+1},\xi):=(L_{2}^{2n+1},\xi_{2})$, since $\pi\left(\sqrt{2}\right)^{2}>\frac{2\pi}{2}$.
@@
We will prove Theorem 1.1 by constructing a contact capacity defined on open subsets of $(L_{k}^{2n+1},\xi_{k})$ and by estimating this contact capacity for the lift of Darboux balls. This contact capacity comes from the spectral selectors defined on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ in [4] by the means of generating functions and Givental’s non-linear Maslov index (see Section 3).
1.2. Non-squeezing in strongly orderable prequantizations
Recall that if $(M,\xi=\ker\alpha)$ is a closed cooriented contact manifold then $(M\times M\times\mathbb{R},\Xi:=\ker(\beta:=\mathrm{pr}_{2}^{*}\alpha-e^{\theta}\mathrm{pr}_{1}^{*}\alpha))$, where $\mathrm{pr}_{i}:M\times M\times\mathbb{R}\to M,\ (x_{1},x_{2},\theta)\mapsto x_{i}$, is a cooriented contact manifold and the diagonal $\Delta:=\{(x,x,0)\ |\ x\in M\}$ is a Legendrian lying inside it. We will call the latter manifold the contact product of $(M,\xi)$222the contact structure on the contact product of $(M,\xi)$ does not depend on the choice of a contact form defining the contact distribution $\xi$ up to isomorphism with itself and will call the $1$-form $\beta$ the product contact form associated to $\alpha$. Using the terminology of [29, 7] we say that $(M,\ker\alpha)$ is strongly orderable if there does not exist any positive contractible loop of Legendrians based at $\Delta$. In particular, if $(M,\ker\alpha)$ is strongly orderable then it is orderable in the sense of [16], i.e. there does not exist any positive contractible loop of contactomorphisms in $\mathrm{Cont}_{0}(M,\ker\alpha)$. See Sections 2.2 and 7.1 for more details about these notions. Colin-Chantraine-Dimitroglou Rizell show in [7, Thm 1.19] that if $(M,\xi)$ is a closed hypertight contact manifold, i.e. there exists a contact form $\alpha$ supporting $\xi$ whose Reeb flow does not have any contractible Reeb orbit, then $(M,\xi)$ is strongly orderable. They also show that any closed contact manifold admitting a Liouville filling whose symplectic homology does not vanish is strongly orderable [7, Thm 1.18].
We say that a closed contact manifold $(M_{1},\xi_{1}^{\prime})$ is a prequantization over a closed symplectic manifold $(W,\omega^{\prime})$ if there exists a $\mathbb{T}^{1}$-principal bundle $\pi_{1}^{\prime}:M_{1}\to W$, where $\mathbb{T}^{1}=\mathbb{R}/2\pi\mathbb{Z}$, and a contact form $\alpha_{1}^{\prime}$ supporting $\xi_{1}^{\prime}$ satisfying ${\pi_{1}^{\prime}}^{*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{1}$ and the Reeb flow of which, is Zoll333all its orbits are periodic with the same minimal period of minimal period $2\pi$, induces the $\mathbb{T}^{1}$-action.
Suppose that $(M_{1},\xi_{1}^{\prime})$ is a closed prequantization over $(W,\omega^{\prime})$, then for any $k\in\mathbb{N}_{>0}$ we denote by $M_{k}:=M_{1}/\mathbb{Z}_{k}$, where $\mathbb{Z}_{k}$ is seen as a subgroup of $\mathbb{T}^{1}$ via the map $[j]\mapsto\frac{2\pi j}{k}$, and $\xi^{\prime}_{k}$ the contact distribution on $M_{k}$ given by the kernel of the unique $1$-form $\alpha^{\prime}_{k}$ whose pull back via the covering $M_{1}\to M_{1}/\mathbb{Z}_{k}$ is $\alpha^{\prime}_{1}$. Finally we denote by $\pi^{\prime}_{k}:M_{k}\to W$ the natural projection. Note that ${\pi^{\prime}_{k}}^{*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{k}$ and that $(M_{k},\xi^{\prime}_{k})$ is a prequantization over $(W,k\omega^{\prime})$ by taking $k\alpha^{\prime}_{k}$ to be the contact form inducing the Zoll Reeb flow of minimal period $2\pi$.
Theorem 1.3.
Suppose that $(M_{1},\xi^{\prime}_{1})$ is a strongly orderable prequantization over a closed symplectic manifold $(W,\omega^{\prime})$ and let $\Psi_{1}:(\mathbb{B}^{2n}(R_{1}),\omega_{0})\hookrightarrow(W,\omega^{\prime})$, $\Psi_{2}:(\mathbb{B}^{2n}(\lambda R_{2}),\omega_{0})\hookrightarrow(W,\omega^{\prime})$ be two symplectic embeddings for some $R_{1}\in(0,\sqrt{2}]$, $R_{2}>0$, $\lambda>\sqrt{3}$, and $a_{1}\in(0,R_{1}],\ a_{2}\in(0,R_{2}]$ be two positive real numbers. For any $k\in\mathbb{N}_{>0}$, if there exists $j\in\mathbb{N}_{>0}$ such that
$\pi a_{1}^{2}>\frac{2\pi}{k}\cdot j\geq\pi a_{2}^{2},$
then $B^{1}_{k}(a_{1})$ cannot be contactly squeezed inside $B^{2}_{k}(a_{2})$, where by definition $B^{i}_{k}(a_{i}):={\pi^{\prime}_{k}}^{-1}(\Psi_{i}(\mathbb{B}^{2n}(a_{i})))\subset(M_{k},\xi^{\prime}_{k})$ for all $i\in\{1,2\}$.
As for Theorem 1.1 we prove Theorem 1.3 by constructing a contact capacity defined on open subsets of $(M_{k},\xi^{\prime}_{k})$ and by estimating this contact capacity for the lift of Darboux balls. The contact capacity this time will come from the order spectral selectors defined in [4] on the universal cover of the isotopy class of the Legendrian $\Delta_{k}$ lying in the contact product of $(M_{k},\xi^{\prime}_{k})$ with itself.
Remark 1.4.
  1. (1)
    In Theorem 1.3 we make the assumption that $R_{1}\leq\sqrt{2}$ so that $\pi R_{1}^{2}$ is not greater than the minimal period of the Reeb flow associated to the contact form $\alpha^{\prime}_{1}$ satisfying ${\pi^{\prime}_{1}}^{*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{1}$ (see the 4th step of the proof of Proposition 8.2 for more details). This technical assumption was not made in Theorem 1.1 because, by Theorem [27, Thm 1] discussed in Remark 1.2, it is automatically satisfied. In particular, when $(M_{1},\xi^{\prime}_{1})$ is a strongly orderable prequantization, one can expect to detect non-squeezing phenomena in $(M_{k},\xi^{\prime}_{k})$ thanks to Theorem 1.3 only for $k>1$.
  2. (2)
    Even if the lens space $(L_{k}^{2n+1},\xi_{k})$ defined in Section 1.1 can be seen as a prequantization over $(\mathbb{C}\mathrm{P}^{n},k\omega=2k\omega_{FS})$, one cannot deduce Theorem 1.1 from Theorem 1.3. First, it is not clear to the author whether contact lens spaces are strongly orderable or not in general. Indeed $(L_{k}^{2n+1},\xi_{k})$, for $n>1$, is not hypertight since it satisfies the Weinstein conjecture (see [22]), i.e. the Reeb flow of any contact form $\alpha$ satisfying $\ker\alpha=\xi_{k}$ admits a periodic orbit, and its fundamental group $\pi_{1}(L_{k}^{2n+1})$ is finite. Moreover, if $n\in\mathbb{N}_{>0}$ is not a power of $2$, Zhou [39] showed that $(L_{2}^{2n-1},\xi_{2})$ is not Liouville fillable. Second, for the reasons discussed in the first point of this Remark 1.4 and the fact that $(L_{1}^{2n+1},\xi_{1})=(\mathbb{S}^{2n+1},\xi_{1})$ is not strongly orderable for $n\geq 1$ (see [15]), Theorem 1.3 cannot detect non-squeezing phenomena in $(L_{p}^{2n+1},\xi_{p})$ when $p$ is a prime number even when it is known that the latter is strongly orderable. In particular, $(L_{2}^{3},\xi_{2})$ is strongly orderable, since the standard symplectic codisk bundle of the $2$-sphere is a Liouville filling whose symplectic homology does not vanish (see Example 1.5 below), but from what we have just said Theorem 1.3 cannot detect non-squeezing phenomena in the latter contact manifold while Theorem 1.1 does (see Remark 1.2).
  3. (3)
    We conjecture that Theorem 1.3 also holds if we assume $(M_{1},\xi^{\prime}_{1})$ to be orderable in the sense of [16], i.e. there does not exist any positive contractible loop of contactomorphisms, which is a priori weaker than to assume $(M_{1},\xi^{\prime}_{1})$ to be strongly orderable. The validity of this conjecture would be an easy consequence of the proof given below in Section 8 together with the validity of the conjecture made in [3] about the spectrality of the (pseudo-)spectral selectors for contactomorphisms.
  4. (4)
    It seems very plausible that with the tools developped in [6] one could derive similar non-squeezing phenomena satisfied by the ideal contact boundary of a semipositive and convex-at-infinity symplectic manifold $W$ whose unit in symplectic homology is not eternal.
Examples 1.5.
  1. (1)
    For any closed manifold $X$, its associated cosphere bundle with its standard contact structure is strongly orderable by [7, Thm. 1.18], since it is Liouville fillable by the standard symplectic codisk bundle and the symplectic homology of the latter does not vanish (see for instance the introduction of [7] or [1, Example 1.6] combined with [34, Thm. 13.3]). If we assume moreover that $X$ admits a Riemaniann metric $g$ so that all its geodesics are closed with length $2\pi$ (see [32, Chap. 3.4] for examples) then the cosphere bundle is, in addition to be strongly orderable, also a prequantization (see [20, Thm. 7.2.5]).
  2. (2)
    If $(W,\omega^{\prime})$ is a closed symplectic manifold so that $[\omega^{\prime}],H_{2}(W,\mathbb{Z})\right>\subset 2\pi\mathbb{Z}$ and $\pi^{\prime}_{1}:M_{1}\to W$ denotes the $\mathbb{R}/2\pi\mathbb{Z}$-principal bundle whose Euler class, after taking its image by the map $H^{2}(W,\mathbb{Z})\to H^{2}(W,\mathbb{R})\simeq H^{2}_{\mathrm{dR}}(W,\mathbb{R})$, is equal to $-[\frac{\omega^{\prime}}{2\pi}]$, then $M_{1}$ carries a contact form $\alpha^{\prime}_{1}$ that turns $(M_{1},\ker\alpha^{\prime}_{1})$ into a prequantization over $(W,\omega^{\prime})$, i.e. ${\pi_{1}^{\prime}}^{*}\omega^{\prime}=d\alpha_{1}^{\prime}$, the Reeb flow of $\alpha_{1}^{\prime}$ is Zoll of minimal period $2\pi$ and it induces the $\mathbb{R}/2\pi\mathbb{Z}$ action (see for instance [20, Chap 7.2]). If we suppose in addition that $(W,\omega^{\prime})$ is symplectically aspherical, i.e. $[\omega^{\prime}],H_{2}^{S}(M,\mathbb{Z})\right>=0$ where $H_{2}^{S}(M,\mathbb{Z})$ is the image of the Hurewitz homomorphism $\pi_{2}(M)\to H_{2}(M,\mathbb{Z})$, then $(M_{1},\xi^{\prime}_{1})$ is hypertight and in particular is strongly orderable thanks to [7, Thm. 1.19]. Therefore we get non-squeezing phenomena thanks to Theorem 1.3 in $(M_{k},\xi^{\prime}_{k})$ for $k>1$ big enough.
@@ 1.3. Discussion
It is interesting to note that the $\alpha_{k}$, and $\alpha^{\prime}_{k}$ Reeb flows on $(L_{k}^{2n+1},\xi_{k})$ and on $(M_{k},\xi^{\prime}_{k})$, as described above in Sections 1.1 and 1.2, are periodic with period $\frac{2\pi}{k}$ for all integers $k\geq 1$. Indeed, the periodicity of the Reeb flow often plays a crucial role in the existence of non trivial global invariants and this article is no exception to this fact (see Lemma 2.2). Furthermore, the global invariants one can deduce from Theorem 1.1 and Theorem 1.3 naturally take value in the discrete set consisting in multiples of this period. Similarly, in the contact manifold $\left(\mathbb{R}^{2n}\times\mathbb{R}/\mathbb{Z},\ker(\alpha_{0}:=\mathrm{d}z-\frac{1}{2}\sum\limits_{i=1}^{n}(x_{i}dy_{i}-y_{i}dx_{i}))\right)$ where the first contact non-squeezing phenomena was described by Eliashberg, Kim and Polterovich in [15], the $\alpha_{0}$-Reeb flow is $1$-periodic, and they show that it is impossible to contactly squeeze $\mathbb{B}^{2n}(A)\times\mathbb{R}/\mathbb{Z}$ inside $\mathbb{B}^{2n}(a)\times\mathbb{R}/\mathbb{Z}$ if there exists an integer $j\in\mathbb{N}_{>0}$ so that $A>j\geq a$.
@@
It seems that another important property that a contact manifold should have so that one could expect meaningful invariant measurements is to be orderable (see 2.2, 7.1 for definitions). For $n>1$ and $k\geq 2$, the sphere $(\mathbb{S}^{2n+1},\ker\alpha_{1})$ and the lens space $(L_{k}^{2n+1},\ker\alpha_{k})$ are perfect representative cases since the sphere is not orderable [15] while the lens space is [22], and, as we have just mentionned, interesting contact measurements exist for the latter but not for the former. In the same direction, a similar observation can be made for bi-invariant metrics on contactomorphisms groups. On the one hand, to the knowledge of the author, the only examples in the litterature of closed contact manifolds $(M,\xi)$ admitting unbounded bi-invariant metrics on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ are the contact manifolds $(M,\xi)$ that are orderable and that admit a periodic Reeb flow [18, 11, 22, 4]. On the other hand, for a large family of closed non-orderable contact manifolds $(M,\xi)$ Courte and Massot show that $\mathrm{Cont}_{0}(M,\xi)$ and $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ have to be bounded [12].
Actually, while convenient but not sufficient, the existence of a periodic Reeb flow is not even a necessary condition to have contact global invariants. For example, the unitary cotangent of the torus $\mathbb{P}^{+}T^{*}\mathbb{T}^{n}$ endowed with its canonical contact structure $\xi_{\mathrm{can}}$ does not admit any periodic Reeb flow when $n>1$ (see for example [8]). But, as briefly mentionned above, one can use an interesting contact global invariant for domains in $(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$, which is called the shape555unlike the other invariant measurements discussed so far, the shape can be used to construct “non-discrete” measurements, to deduce non-squeezing phenomena and other rigidity results concerning this contact manifold [14, 5, 33, 16]. It would therefore be interesting to investigate whether the groups $\widetilde{\mathrm{Cont}_{0}}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ and $\ \mathrm{Cont}_{0}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ are bounded or not (see also [3, Prop. 4.14, Rem. 4.15]).
@@
In Section 2 we set some conventions and give basic definitions and properties of orderability and spectral selectors. In Section 3 we briefly recall the construction and main properties of the spectral selectors on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ constructed in [4]. In Section 4 we define the above mentionned contact capacity for open domains of $(L_{k}^{2n+1},\xi_{k})$ coming from the spectral selectors. In this same Section 4 we state - without prooving - some properties of this contact capacity (see Theorem 4.1) which allow us to give the proof of Theorem 1.1. In Section 5 we show a capacity-energy inequality satisfied by the contact capacity. In Section 6 we prove Theorem 4.1 by estimating and computing the contact capacity of domains that are lifts of Darboux balls to $(L_{k}^{2n+1},\xi_{k})$ using the capacity-energy inequality of Section 5 and some explicit constructions. In Section 7 we define the spectral selectors on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ when $(M,\xi)$ is a strongly orderable closed contact manifold. Finally in Section 8 we use these spectral selectors to define a contact capacity for domains in strongly orderable prequantizations and prove Theorem 1.3.